# Stable non-Hausdorff Riemannian foliations

### v2

@@ 1. Introduction
Recent work by the author and Zeiser [6] aimed to find examples of stable foliations other than the Hausdorff foliations considered above. To explain the main result of [6], recall that for any foliation $\mathcal{F}$ there is a canonical representation $\nabla$ of the Lie algebroid $T\mathcal{F}$ on the normal bundle $N\mathcal{F}$, called Bott connection. Hence, one obtains a complex $\big{(}\Omega^{\bullet}(\mathcal{F},N\mathcal{F}),d_{\nabla}\big{)}$ with cohomology groups $H^{\bullet}(\mathcal{F},N\mathcal{F})$. Heitsch [8] showed that first order deformations of the foliation $\mathcal{F}$, modulo those obtained by applying an isotopy to $\mathcal{F}$, are given by elements in the cohomology group $H^{1}(\mathcal{F},N\mathcal{F})$. Therefore, vanishing of $H^{1}(\mathcal{F},N\mathcal{F})$ can be interpreted heuristically as the infinitesimal requirement for stability of $\mathcal{F}$. The main result in [6] states that the vanishing of $H^{1}(\mathcal{F},N\mathcal{F})$ actually implies that $\mathcal{F}$ is stable, at least when $\mathcal{F}$ is Riemannian. Recall that $\mathcal{F}$ is Riemannian if $M$ admits a metric that is bundle-like with respect to $\mathcal{F}$ [16]. Equivalently, $\mathcal{F}$ is locally defined by Riemannian submersions.
@@
On a compact manifold $M$, Hausdorff foliations are exactly the Riemannian foliations with all leaves compact. Moreover, if the generic leaf $L$ of a Hausdorff foliation $\mathcal{F}$ satisfies $H^{1}(L)=0$, then also the cohomology group $H^{1}(\mathcal{F},N\mathcal{F})$ vanishes [7]. So our Theorem 1.2 implies Theorem 1.1 by Hamilton. However, it was still unclear whether Theorem 1.2 admits new examples that are not already covered by Theorem 1.1. The present note aims to fill this gap. We construct examples of Riemannian foliations $\mathcal{F}$ with dense leaves for which $H^{1}(\mathcal{F},N\mathcal{F})$ vanishes, showing that Theorem 1.2 is indeed more general than Theorem 1.1.
@@
The proof of the Main Theorem relies on recent work by El Kacimi Alaoui [3] which computes the cohomology group $H^{1}(\mathcal{F},N\mathcal{F})$ for developable foliations $\mathcal{F}$. These are foliations $\mathcal{F}$ whose lift $\widehat{\mathcal{F}}$ to some covering $\widehat{M}$ of $M$ is given by a fibration. Note that the suspension foliation $\mathcal{F}$ of a homomorphism $\varphi:\pi_{1}(B)\rightarrow G$ is of this type. The main result of [3] shows that $H^{1}(\mathcal{F},N\mathcal{F})$ is given by the first group cohomology $H^{1}\left(\pi_{1}(B),\mathfrak{X}(G)\right)$. Here the discrete group $\pi_{1}(B)$ acts on $G$ by left translations (through the homomorphism $\varphi$) and therefore on $\mathfrak{X}(G)$ by pushforward. Endowing $G$ with a left invariant metric, the representation of $\pi_{1}(B)$ on $\mathfrak{X}(G)$ is orthogonal for the induced $L^{2}$-inner product. The assumption that $\pi_{1}(B)$ has Kazhdan’s property (T) ensures that its cohomology $H^{1}\left(\pi_{1}(B),V\right)$ vanishes whenever $V$ is a real Hilbert space and the representation of $\pi_{1}(B)$ on $V$ is orthogonal [1, Chapter 2]. This does not immediately imply the vanishing of $H^{1}\left(\pi_{1}(B),\mathfrak{X}(G)\right)$ but rather that of $H^{1}\left(\pi_{1}(B),L^{2}\mathfrak{X}(G)\right)$, where $L^{2}\mathfrak{X}(G)$ is the Hilbert space of square integrable vector fields on $G$ for the $L^{2}$-inner product. However, the fact that $\varphi(\pi_{1}(B))\subset G$ is dense implies that the action of $\pi_{1}(B)$ on $G$ is ergodic, hence a regularity result by Lubotzky-Zimmer [11] can be applied. It shows that a $1$-cocycle $\pi_{1}(B)\rightarrow\mathfrak{X}(G)$, which a priori has a primitive defined in terms of an element in $L^{2}\mathfrak{X}(G)$, admits a primitive coming from a smooth vector field. Consequently, the first group cohomology $H^{1}\left(\pi_{1}(B),\mathfrak{X}(G)\right)$ vanishes.
@@ 2. Suspension foliations
the associated suspension $(B_{\varphi},\mathcal{F})$ is a Riemannian foliation (in fact, a Lie $G$-foliation). We will only consider homomorphisms $\varphi$ for which the image $H:=\varphi(\pi_{1}(B))$ is a dense subgroup of $G$. In that case, the suspension foliation $\mathcal{F}$ on $B_{\varphi}$ is a Lie $G$-foliation with dense leaves. This follows immediately from the fact that $\mathcal{F}$ is a transversely parallelizable foliation on a compact connected manifold, all of whose basic functions are constant [13, Thm. 4.24]. To see this, note that a frame of left invariant vector fields on $G$ gives a transverse parallelism for the horizontal foliation on $\widehat{B}\times G$, which descends to a transverse parallelism for $\mathcal{F}$ on $B_{\varphi}$. Moreover, the $\mathcal{F}$-basic functions on $B_{\varphi}$ are constant since
@@ 3.1. Infinitesimal deformations of developable foliations
The work [8] of Heitsch shows that infinitesimal deformations of the foliation $\mathcal{F}$ are one-cocycles in $\big{(}\Omega^{\bullet}(\mathcal{F};N\mathcal{F}),d_{\nabla}\big{)}$. Moreover, if a smooth deformation of $\mathcal{F}$ is obtained applying an isotopy to $\mathcal{F}$, then the corresponding infinitesimal deformation is a one-coboundary in $\big{(}\Omega^{\bullet}(\mathcal{F};N\mathcal{F}),d_{\nabla}\big{)}$. Consequently, the vanishing of the cohomology group $H^{1}(\mathcal{F},N\mathcal{F})$ can be interpreted heuristically as the infinitesimal requirement for stability of the foliation $\mathcal{F}$. In case the foliation $\mathcal{F}$ is developable, then $H^{1}(\mathcal{F},N\mathcal{F})$ can be expressed in terms of group cohomology, as shown by El Kacimi Alaoui [3]. We now recall this result.
@@
Let $\varphi:\pi_{1}(B)\rightarrow\text{Diff}(T)$ be a group homomorphism, where $B$ and $T$ are compact connected manifolds. The suspension foliation $\mathcal{F}$ on $B_{\varphi}$ lifts under $\pi:\widehat{B}\times T\rightarrow B_{\varphi}$ to the foliation by fibers of $\widehat{B}\times T\rightarrow T$. Since the universal cover $\widehat{B}$ is simply connected, we have $H^{1}(\widehat{B})=0$. Hence, Proposition 3.3 implies that
$H^{1}(\mathcal{F},N\mathcal{F})\cong H^{1}\big{(}\pi_{1}(B),\mathfrak{X}(T)\big{)}.$
@@ 3.2. Actions of Kazhdan groups
Kazhdan’s property (T) is relevant for our purpose because of the following: a countable discrete group $\Gamma$ has property (T) exactly when the first group cohomology $H^{1}(\Gamma,V)$ vanishes for every orthogonal representation of $\Gamma$ on a real Hilbert space $V$ (see [1, Chapter 2]). This statement is not directly applicable in our situation because the representations that we encounter are no Hilbert spaces, see (2). To overcome this issue, we need to invoke the following result by Lubotzky-Zimmer [11, Theorem 4.1]. We only state a particular case of their result, which holds under more general assumptions that are implied by property (T).
@@ 3.3. Proof of Main Theorem
The action of $\pi_{1}(B)$ on $G$ is isometric with respect to any left invariant Riemannian metric on $G$. Moreover, it is well-known that if $\Gamma\subset G$ is a dense subgroup of a compact Lie group, then the action by left translations $\Gamma\curvearrowright G$ is ergodic with respect to the normalized Haar measure on $G$ [15, Lemma 2.4.1]. Since by assumption $\varphi(\pi_{1}(B))\subset G$ is dense, it follows that the action $\pi_{1}(B)\curvearrowright(G,\mu)$ is ergodic. Moreover, the fundamental group $\pi_{1}(B)$ of a compact manifold $B$ is finitely generated. Applying Prop. 3.7, we see that the cohomology group (3) vanishes. Stability of $\mathcal{F}$ immediately follows from Thm. 1.2 in the Introduction. ∎
@@ 4. An example
The aim of this section is to construct a discrete group $\Gamma$ which can serve as the fundamental group $\pi_{1}(B)$ in Thm. 3.8. We will obtain $\Gamma$ as a suitable arithmetic lattice in $SO(3,2)$ which can be realized as a dense subgroup of $SO(5)$. The lattice in question was considered before in connection with the Ruziewicz problem, which concerns rotationally invariant measures on spheres [10, §3.4].
@@ 4.1. Arithmetic lattices
We start by recalling an arithmetic construction of lattices, following [12, Chapter IX, §1.7]. Fix an integer $n\geq 3$ and a subfield $K\subset\mathbb{C}$ which is a finite extension of $\mathbb{Q}$. Denote by $d=[K:\mathbb{Q}]$ the degree of the extension. Assume that
@@
[12, Chapter IX, §1.7] Assume that $\mathcal{T}\neq\mathcal{R}$ and that the group $SO_{\Phi}$ is almost $K$-simple. Then $SO_{\Phi}(\mathcal{O})$ is an arithmetic lattice in $\prod_{\sigma\in\mathcal{S}}SO_{\Phi_{\sigma}}(k_{\sigma})$.
@@ 4.2. An example
  • Lattices in connected semisimple Lie groups are finitely presented [5, §1.1]. Viewing $\Gamma$ as a lattice inside $SO(3,2)$, this implies that $\Gamma$ is finitely presented, even though $SO(3,2)$ has $2$ connected components. To see this, let us denote by $SO(3,2)^{+}$ its identity component, which is a normal index $2$ subgroup $SO(3,2)^{+}\subset SO(3,2)$. The fact that $SO(3,2)^{+}$ is an open subgroup implies that $\Gamma\cap SO(3,2)^{+}$ is a lattice inside $SO(3,2)^{+}$, hence $\Gamma\cap SO(3,2)^{+}$ is finitely presented by the above. Moreover, since
@@
  • We have the compact, connected Lie group $G:=SO(5)$ and a group homomorphism $\varphi:\Gamma\rightarrow SO(5)$ given by the restriction of the second projection in $SO(3,2)\times SO(5)$. It was shown in [10, Prop. 3.4.3] that the projection of $\Gamma$ in $SO(5)$ is dense.
@@
We obtain an arithmetic lattice $\Gamma\subset SO(n-2,2)\times SO(n)$ with property (T), which projects to a dense subgroup in $SO(n)$ [10, Prop. 3.4.3].
The method fails for $n<5$. In fact, Thm. 3.8 admits no examples in which $G=SO(n)$ with $n<5$. Indeed, if $\varphi:\Gamma\rightarrow SO(n)$ is a group homomorphism where $\Gamma$ is a discrete Kazhdan group and $n<5$, then the image $\varphi(\Gamma)$ is finite. This was shown by Zimmer in [17, Thm. 7] for $SO(3)$ and $SO(4)$. It is immediate in the case of $SO(2)$, because of the following. Since $SO(2)$ is abelian, the group homomorphism $\varphi:\Gamma\rightarrow SO(2)$ factors through the abelianization Γ/[$\Gamma/[\Gamma,\Gamma]$. Because the abelianization of a discrete group with property (T) is finite [1, Cor. 1.3.6], it follows that the image $\varphi(\Gamma)\subset SO(2)$ is finite. This shows that the suspension foliation of any homomorphism $\varphi:\Gamma\rightarrow SO(n)$, with $\Gamma$ a discrete Kazhdan group and $n<5$, is Hausdorff since all of its leaves are compact (see §2).

### v3

@@ 1. Introduction
Aside from the Hausdorff foliations appearing in Thm. 1.1, very few examples of stable foliations are known. El Kacimi Alaoui and Nicolau [4] established stability for a particular class of non-Hausdorff foliations obtained by suspending linear foliations on tori by suitable linear Anosov diffeomorphisms. Their proof is an ad hoc application of the method used by Hamilton to prove Thm. 1.1.
Recent work by the author and Zeiser [7] aimed to establish a stability result for foliations that extends Thm. 1.1. To explain the main result of [7], recall that for any foliation $\mathcal{F}$ there is a canonical representation $\nabla$ of the Lie algebroid $T\mathcal{F}$ on the normal bundle $N\mathcal{F}$, called Bott connection. Hence, one obtains a complex $\big{(}\Omega^{\bullet}(\mathcal{F},N\mathcal{F}),d_{\nabla}\big{)}$ with cohomology groups $H^{\bullet}(\mathcal{F},N\mathcal{F})$. Heitsch [9] showed that first order deformations of $\mathcal{F}$, modulo those obtained by applying an isotopy to $\mathcal{F}$, are given by elements in $H^{1}(\mathcal{F},N\mathcal{F})$. Therefore, vanishing of $H^{1}(\mathcal{F},N\mathcal{F})$ can be interpreted heuristically as the infinitesimal requirement for stability of $\mathcal{F}$. The main result in [7] states that vanishing of $H^{1}(\mathcal{F},N\mathcal{F})$ actually implies that $\mathcal{F}$ is stable, at least when the foliation $\mathcal{F}$ is Riemannian. Recall here that $\mathcal{F}$ is Riemannian if the manifold $M$ admits a Riemannian metric that is bundle-like with respect to $\mathcal{F}$ [17]. Equivalently, the foliation $\mathcal{F}$ is locally defined by Riemannian submersions.
@@
To relate this result with Thm. 1.1, note that on a compact manifold $M$, Hausdorff foliations are exactly the Riemannian foliations with all leaves compact. Moreover, if the generic leaf $L$ of a Hausdorff foliation $\mathcal{F}$ satisfies $H^{1}(L)=0$, then also the cohomology group $H^{1}(\mathcal{F},N\mathcal{F})$ vanishes [8]. So Thm. 1.2 implies Thm. 1.1. However, it was still unclear whether Thm. 1.2 admits new examples that are not already covered by Thm. 1.1. This note aims to fill that gap. We construct examples of Riemannian foliations $\mathcal{F}$ with dense leaves for which $H^{1}(\mathcal{F},N\mathcal{F})$ vanishes, showing that Thm. 1.2 is indeed more general than Thm. 1.1. To our knowledge, these are the first examples of stable Riemannian foliations that are not Hausdorff – note that the stable foliations in [4] are not Riemannian.
@@
The proof of the Main Theorem relies on recent work by El Kacimi Alaoui [3] which computes the cohomology group $H^{1}(\mathcal{F},N\mathcal{F})$ for developable foliations $\mathcal{F}$. These are foliations $\mathcal{F}$ whose lift $\widehat{\mathcal{F}}$ to some covering $\widehat{M}$ of $M$ is given by a fibration. Note that the suspension foliation $\mathcal{F}$ of a homomorphism $\varphi:\pi_{1}(B)\rightarrow G$ is of this type. The main result of [3] shows that $H^{1}(\mathcal{F},N\mathcal{F})$ is given by the first group cohomology $H^{1}\left(\pi_{1}(B),\mathfrak{X}(G)\right)$. Here the discrete group $\pi_{1}(B)$ acts on $G$ by left translations (through the homomorphism $\varphi$) and therefore on $\mathfrak{X}(G)$ by pushforward. Endowing $G$ with a left invariant metric, the representation of $\pi_{1}(B)$ on $\mathfrak{X}(G)$ is orthogonal for the induced $L^{2}$-inner product. The assumption that $\pi_{1}(B)$ has Kazhdan’s property (T) ensures that the cohomology group $H^{1}\left(\pi_{1}(B),V\right)$ vanishes whenever $V$ is a real Hilbert space and the representation of $\pi_{1}(B)$ on $V$ is orthogonal [1, Chapter 2]. This does not immediately imply the vanishing of $H^{1}\left(\pi_{1}(B),\mathfrak{X}(G)\right)$ but rather that of $H^{1}\left(\pi_{1}(B),L^{2}\mathfrak{X}(G)\right)$, where $L^{2}\mathfrak{X}(G)$ is the Hilbert space of square integrable vector fields on $G$ for the $L^{2}$-inner product. However, the fact that $\varphi(\pi_{1}(B))\subset G$ is dense implies that the action of $\pi_{1}(B)$ on $G$ is ergodic, hence a regularity result by Lubotzky-Zimmer [12] can be applied. It shows that a $1$-cocycle $\pi_{1}(B)\rightarrow\mathfrak{X}(G)$, which a priori has a primitive defined in terms of an element in $L^{2}\mathfrak{X}(G)$, admits a primitive coming from a smooth vector field. Consequently, the first group cohomology $H^{1}\left(\pi_{1}(B),\mathfrak{X}(G)\right)$ vanishes.
@@ 2. Suspension foliations
the associated suspension $(B_{\varphi},\mathcal{F})$ is a Riemannian foliation (in fact, a Lie $G$-foliation). We will only consider homomorphisms $\varphi$ for which the image $H:=\varphi(\pi_{1}(B))$ is a dense subgroup of $G$. In that case, the suspension foliation $\mathcal{F}$ on $B_{\varphi}$ is a Lie $G$-foliation with dense leaves. This follows immediately from the fact that $\mathcal{F}$ is a transversely parallelizable foliation on a compact connected manifold, all of whose basic functions are constant [14, Thm. 4.24]. To see this, note that a frame of left invariant vector fields on $G$ gives a transverse parallelism for the horizontal foliation on $\widehat{B}\times G$, which descends to a transverse parallelism for $\mathcal{F}$ on $B_{\varphi}$. Moreover, the $\mathcal{F}$-basic functions on $B_{\varphi}$ are constant since
@@ 3.1. Infinitesimal deformations of developable foliations
Heitsch [9] showed that infinitesimal deformations of the foliation $\mathcal{F}$ are one-cocycles in the complex $\big{(}\Omega^{\bullet}(\mathcal{F};N\mathcal{F}),d_{\nabla}\big{)}$. Moreover, if a smooth deformation of $\mathcal{F}$ is obtained applying an isotopy to $\mathcal{F}$, then the corresponding infinitesimal deformation is a one-coboundary in $\big{(}\Omega^{\bullet}(\mathcal{F};N\mathcal{F}),d_{\nabla}\big{)}$. Hence, vanishing of the cohomology group $H^{1}(\mathcal{F},N\mathcal{F})$ can be interpreted heuristically as the infinitesimal requirement for stability of the foliation $\mathcal{F}$. In case the foliation $\mathcal{F}$ is developable, then $H^{1}(\mathcal{F},N\mathcal{F})$ can be expressed in terms of group cohomology, as shown by El Kacimi Alaoui [3]. We now recall this result.
@@
Let $\varphi:\pi_{1}(B)\rightarrow\text{Diff}(T)$ be a group homomorphism, where $B$ and $T$ are compact connected manifolds. The suspension foliation $\mathcal{F}$ on $B_{\varphi}$ lifts under $\pi:\widehat{B}\times T\rightarrow B_{\varphi}$ to the foliation by fibers of $\widehat{B}\times T\rightarrow T$. Since the universal cover $\widehat{B}$ is simply connected, we have $H^{1}(\widehat{B})=0$. Hence, Prop. 3.3 implies that $H^{1}(\mathcal{F},N\mathcal{F})\cong H^{1}\big{(}\pi_{1}(B),\mathfrak{X}(T)\big{)}.$
@@ 3.2. Actions of Kazhdan groups
Kazhdan’s property (T) is relevant for our purpose because of the following: a countable discrete group $\Gamma$ has property (T) exactly when the first group cohomology $H^{1}(\Gamma,V)$ vanishes for every orthogonal representation of $\Gamma$ on a real Hilbert space $V$ [1, Chapter 2]. This statement is not directly applicable in our situation because the representations that we encounter are no Hilbert spaces, see (2). To overcome this issue, we need to invoke the following result by Lubotzky-Zimmer [12, Thm. 4.1]. We only state a particular case of their result, which holds under more general assumptions that are implied by property (T).
@@ 3.3. Proof of Main Theorem
The action of $\pi_{1}(B)$ on $G$ is isometric with respect to any left invariant Riemannian metric on $G$. Moreover, it is well-known that if $\Gamma\subset G$ is a dense subgroup of a compact Lie group, then the action by left translations $\Gamma\curvearrowright G$ is ergodic with respect to the normalized Haar measure on $G$ [16, Lemma 2.4.1]. Since by assumption $\varphi(\pi_{1}(B))\subset G$ is dense, it follows that the action $\pi_{1}(B)\curvearrowright(G,\mu)$ is ergodic. Moreover, the fundamental group $\pi_{1}(B)$ of a compact manifold $B$ is finitely generated. Applying Prop. 3.7, we see that the cohomology group (3) vanishes. Stability of $\mathcal{F}$ immediately follows from Thm. 1.2 in the Introduction. ∎
@@ 4. An example
The aim of this section is to construct a discrete group $\Gamma$ which can serve as the fundamental group $\pi_{1}(B)$ in Thm. 3.8. We will obtain $\Gamma$ as a suitable arithmetic lattice in $SO(3,2)$ which can be realized as a dense subgroup of $SO(5)$. The lattice in question was considered before in connection with the Ruziewicz problem, which concerns rotationally invariant measures on spheres [11, §3.4].
@@ 4.1. Arithmetic lattices
We start by recalling an arithmetic construction of lattices, following [13, Chapter IX, §1.7]. Fix an integer $n\geq 3$ and a subfield $K\subset\mathbb{C}$ which is a finite extension of $\mathbb{Q}$. Denote by $d=[K:\mathbb{Q}]$ the degree of the extension. Assume that
@@
[13, Chapter IX, §1.7] Assume that $\mathcal{T}\neq\mathcal{R}$ and that the group $SO_{\Phi}$ is almost $K$-simple. Then $SO_{\Phi}(\mathcal{O})$ is an arithmetic lattice in $\prod_{\sigma\in\mathcal{S}}SO_{\Phi_{\sigma}}(k_{\sigma})$.
@@ 4.2. An example
  • Lattices in connected semisimple Lie groups are finitely presented [6, §1.1]. Viewing $\Gamma$ as a lattice inside $SO(3,2)$, this implies that $\Gamma$ is finitely presented, even though $SO(3,2)$ has $2$ connected components. To see this, let us denote by $SO(3,2)^{+}$ its identity component, which is a normal index $2$ subgroup $SO(3,2)^{+}\subset SO(3,2)$. The fact that $SO(3,2)^{+}$ is an open subgroup implies that $\Gamma\cap SO(3,2)^{+}$ is a lattice inside $SO(3,2)^{+}$, hence $\Gamma\cap SO(3,2)^{+}$ is finitely presented by the above. Moreover, since
@@
  • We have the compact, connected Lie group $G:=SO(5)$ and a group homomorphism $\varphi:\Gamma\rightarrow SO(5)$ given by the restriction of the second projection in $SO(3,2)\times SO(5)$. It was shown in [11, Prop. 3.4.3] that the projection of $\Gamma$ in $SO(5)$ is dense.
@@
We obtain an arithmetic lattice $\Gamma\subset SO(n-2,2)\times SO(n)$ with property (T), which projects to a dense subgroup in $SO(n)$ [11, Prop. 3.4.3].
The method fails for $n<5$. In fact, Thm. 3.8 admits no examples in which $G=SO(n)$ with $n<5$. Indeed, if $\varphi:\Gamma\rightarrow SO(n)$ is a group homomorphism where $\Gamma$ is a discrete Kazhdan group and $n<5$, then the image $\varphi(\Gamma)$ is finite. This was shown by Zimmer in [18, Thm. 7] for $SO(3)$ and $SO(4)$. It is immediate in the case of $SO(2)$, because of the following. Since $SO(2)$ is abelian, the group homomorphism $\varphi:\Gamma\rightarrow SO(2)$ factors through the abelianization Γ/[$\Gamma/[\Gamma,\Gamma]$. Because the abelianization of a discrete group with property (T) is finite [1, Cor. 1.3.6], it follows that the image $\varphi(\Gamma)\subset SO(2)$ is finite. This shows that the suspension foliation of any homomorphism $\varphi:\Gamma\rightarrow SO(n)$, with $\Gamma$ a discrete Kazhdan group and $n<5$, is Hausdorff since all of its leaves are compact (see §2).